The widths of two similar rectangles are 25 ft. and 20 ft. What is the ratio of the perimeters? Of the ares?
Answer by htmentor(1343)

Corresponding sides of similar rectangles have the same ratio. The ratio of the widths = 20/25 = 4/5
This means that the ratio of the lengths is also 4/5.
Let x = the longer length
Then the shorter length = (4/5)x
The perimeters of the two rectangles are P1 = 2*20 + 2*(4/5)x and P2 = 2*25 + 2x
P1 = 4(10 + (2/5)x)
P2 = 5(10 + (2/5)x)
The ratio P1/P2 = 4/5
The areas are A1 = 20*(4/5)x and A2 = 25x
The ratio A1/A2 = 16x/25x = (4/5)^2 = 0.64